# Strong primordial inhomogeneities induced by axion-like scalar field

### v2

@@ 2 Thin shell approximation
Let us denote wall’s surface as $\Sigma$, which is 3-dimensional spacelike or timelike (this would affect $\epsilon$, it would be $+1$ [spacelike] or $-1$ [timelike]) hyper-surface embedded in a 4-dimensional spacetime ($\mathcal{M},g_{\mu\nu}$). We put subscripts $\pm$ to denote the value on each side of the hypersurface $\Sigma$. Following aforementioned papers:
@@
    $[K_{\mu\nu}]^{\pm}=8\pi\epsilon\left(-S_{\mu\nu}+\cfrac{1}{2}\,Sh_{\mu\nu}\right).$ (7)
@@
if one assume the smooth crossing of the wall by fluid ($[u_{\mu}\xi^{\mu}]^{\pm}=0$).
@@ 3 Axion-like field in Kantowski-Sachs metric
Figure 1: The plot of the solution of (36) for different values of of $l$, where the interesting observation is the impact of $l$ changing slowly for $l<-2$.
@@
is plotted in Fig.(2) 222 Note that the parameter has also been scaled appropriately with regards to the differential equation with $\theta$ variable.
Figure 2: The plot of the equation of state parameter for the scalar field. We see that scalar field could behave differently depending on the value of $l$ and in particular, if $l=-1.5$ then scalar field behaves like non-relativistic matter.
@@ Acknowledgements
We are grateful to Yu.N. Eroshenko for fruitful discussions. The research by M.Yu.K. was carried out in the Southern Federal University with financial support of the Ministry of Science and Higher Education of the Russian Federation (State contract GZ0110/23-10-IF).

### v3

@@ 2 Thin shell approximation
Let us denote wall’s surface as $\Sigma$, which is 3-dimensional spacelike or timelike (this would affect $\epsilon$, it would be +1+1 [spacelike] or 1-1 [timelike]) hyper-surface embedded in a 4-dimensional spacetime ($\mathcal{M},g_{\mu\nu}$). We put subscripts $\pm$ to denote the value on each side of the hypersurface $\Sigma$. Following aforementioned papers:
@@
    $K_{\mu\nu}]^{\pm}=8\pi\epsilon\left(-S_{\mu\nu}+\cfrac{1}{2}\,Sh_{\mu\nu}\right).$ (7)
@@
if one assume the smooth crossing of the wall by fluid ($u_{\mu}\xi^{\mu}]^{\pm}=0$).
@@ 3 Axion-like field in Kantowski-Sachs metric
Figure 1: The plot of the solution of (36) for different values of of $l$, where the interesting observation is the impact of $l$ changing slowly for l<2l<-2.
@@
is plotted in Fig.(2) [^2] Note that the parameter has also been scaled appropriately with regards to the differential equation with $\theta$ variable.
Figure 2: The plot of the equation of state parameter for the scalar field. We see that scalar field could behave differently depending on the value of $l$ and in particular, if l=1.5l=-1.5 then scalar field behaves like non-relativistic matter.
@@ Acknowledgements
We are grateful to Yu.N. Eroshenko for fruitful discussions.